Evaluate tan 30° without using a calculator by using ratios in a reference triangle.


help me please show work please i did it it gives me -5.41
1
Expert's answer
2014-05-26T05:09:17-0400
As tan(x)=sin(x)/cos(x) we have that tan(30°)=sin(30°)/cos(30°)=1/2:(3)^(1/2)/2=1/(square root of 3).